News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 27 October 2000 11.00am EDT

During last week the observing schedule was completed as planned with the exception of approximately 6 hours of ACIS observing that was impacted when the ACIS Digital Processing Assembly (DPA-A) powered off at 11:42 am EDT on 26 Sept. The turn-off was detected during a real-time pass and following engineering assessment, the DPA was powered on and ACIS returned to science operation. ACIS has operated nominally since then and the cause of the turn-off is being investigated by the ACIS engineering team. Observations of HD100546, HD104237 and a portion of Sgr A were impacted and will be rescheduled.

Real-time procedures were completed during the week to recover the data stored on SSR-A that were lost on the ground (20 Oct), retract the LETG as planned (25 Oct), load the software patches and disable an Electrical Power System safing monitor in preparation for the lunar eclipse on 27 Oct (25 Oct), and to power-on the ACIS DPA and resume ACIS science operation (26 Oct). A routine check of the CPE Warm Start Count was also completed on 23 Oct and showed the value remained at 1.

The planned observations for the coming week are as follows.

------------------------------------------
SN 1999EM               ACIS-S       Oct 29 
GSGP4X:082              ACIS-S       Oct 30 
GSGP4X:057              ACIS-S
GSGP4X:069              ACIS-S
Radiation Belt
W44                     ACIS-S       Oct 31 
V1701+6414              ACIS-I
M31                     ACIS-S       Nov 1  
NGC 891                 ACIS-S
Radiation Belt                       Nov 2  
NGC 3623                ACIS-S       Nov 3  
PG1115+080              ACIS-S
Crab Nebula             ACIS-S
A401                    ACIS-I
NGC 1700                ACIS-S
3C 207                  ACIS-S       Nov 4  
AGX J104645+1140        ACIS-I
AGX J111432+4056        ACIS-I
Radiation Belt
M31                     ACIS-S       Nov 5

------------------------------------------

With the above noted, the spacecraft subsystems continued to operate nominally.

Return to Status Reports